answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the chance of drawing a number7 of cards?

The probability of drawing a seven in a standard deck of 52 cards is 4 in 52, or 1 in 13, or about 0.07692.

What skills do you need to be a kindergarten teacher?

well, you have to be patient, thoughtful, kind, and you gotta reduce your temper

1896 Indian head penny?

$2 to $26,500 depending upon the condition of the coin.

A card drawn from 52 cards the probability of getting a face card or spade?

The probability of drawing a spade is 1/4 (13/52).

The probability of drawing a face card is 3/13 (12/52).

The probability of drawing a spade face card is 3/52 (3/52).

13 + 12 - 3 = 22

The probability of drawing a spade or a face card is 11/26 (22/52).

If you draw a card from a standard deck of cards what is the probability of drawing a nine?

There are 4 nines in a 52 card deck so the probability of drawing a nine is 4/52 or 1/13.

Which is 7%

What is the probability of getting a 1 or 2 on a six sided dice?

The probability of landing on any 1 given number is 1:6. Specific numbers don't matter. It is all odds.

Just multiply by two so that we get a 2:6 probability that you will land on a 1 or 2. this reduces to 1:3 or 33.3%

What is the probability of tossing the dice and it landing on a 2 and then tossing it a second time and it landing of a number greater than 4?

First probability of getting 2 is one in six: 1/6

probability of getting no greater than 4 is two chances in six (must be a 5 or 6): 2/6=1/3

so probability = 1/6*1/3= 1/18.

What is the probability of a homozygous recessive offspring?

It depends. If it's a heterozygous cross, (Tt x Tt), there's a 25% chance. If it's a homozygous dominant cross (TT x TT), the chance is 0%. Neither parent has the alleles for a recessive trait, so none of their offspring can have the recessive trait. If it's a homozygous recessive cross (tt x tt), there's a 100% chance. The only alleles the parents can pass on are recessive.

What is the probability of Mendel's principle?

The probability is 1. Mendel's principle exists.

How do you combine probabilities?

like for or we use + and for 'and' we use multiplication

In a family of 5 what is the probability that you will have 3 girls (please show your work)?

The probability of a female is approx 0.4831 across the world. If you assume (rather dubious) that the genders of children in a family are independent then the probability is approx 0.3012


In a family of 5, the probability of 3 girls and 2 boys is

5C3*(0.4831)3*(0.5169)2 = 0.3012

What is the interpretation of probability?

It is a measure of the likelihood of a specified event occurring in an experiment.

What is the probability of getting a seven on either of two rolls with a nine sided die?

It's 1/9 on the first roll and 1/9 on the second roll, so you have

a 2/9 = 22.22% chance of success on either one or the other.

What is the probability of flipping a coin four times and getting the same side all four times?

Each flip has two possible outcomes and they are independent events, so there are 24 = 16 possible results.

Of these, only 2 (HHHH, TTTT) are the same 4 each time, Thus:

probability = 2/16 = 1/8

What are the chances of having a girl after four boys in a row?

Just about 50/50. The sex of earlier children has no bearing on the sex of the next child to be born. Each pregnancy has a (just about) even chance of being a boy or a girl.